# Supernova Electron-Neutrino Interactions with Xenon in the nEXO Detector

### v1

@@ VII Acknowledgments
We would like to thank the authors of Ref. [35] for providing their calculations of the GT strength used in this publication, and Steven Gardiner for his continued development and support of the MARLEY event generator. The authors gratefully acknowledge support for nEXO from the Office of Nuclear Physics within DOE’s Office of Science, and NSF in the United States; from NSERC, CFI, FRQNT, NRC, and the McDonald Institute (CFREF) in Canada; from IBS in Korea; and from CAS and NSFC in China. This work was supported in part by Laboratory Directed Research and Development (LDRD) programs at Brookhaven National Laboratory (BNL), Lawrence Livermore National Laboratory (LLNL), Oak Ridge National Laboratory (ORNL), Pacific Northwest National Laboratory (PNNL), and SLAC National Accelerator Laboratory. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344.
Appendix A Individual output channel energy distributions
@@
Figs.  10, 11, and 12 show the fraction of the incident neutrino energy transferred to the various output channels for CC $\nu_{e}-{}^{136}$eCC, $\nu_{e}-{}^{136}$eCC, and neutrino-electron elastic scattering events.
Appendix B Individual transfer matrices
@@
Figs. 13, 14, and 15 show the transfer matrices for each of the scattering components described within the text, generated with one-million simulated and reconstructed events with a uniform neutrino distribution from 0 to 100 MeV. These are combined with isotopic and cross section weighting in Fig. 6 of the main text.
Appendix C ${}^{132}\mbox{Xe}$ inclusive cross section comparison
As a further test of MARLEY, we calculate predictions for the $\nu_{e}-{}^{132}$eCC cross section as a function of energy using GT strength from Ref. [35] and an IAS location predicted by Eq. 2. Fig. 16 compares MARLEY’s predictions with existing calculations for 132Xe from Refs. [21, 23, 22]. Refs. [21, 22] use independent theoretical calculations that include forbidden transitions, and use an approach to the modeling the Coulomb correction similar to that used by MARLEY. Ref. [21] assumes an unquenched $g_{A}$ of $\sim$1.26. Ref. [23] uses the same theoretical GT matrix elements as we use in MARLEY, but assumes a different location for the IAS in 132Cs, an unquenched $g_{A}$ of $\sim$1.26, and uses a different approach to modeling the Coulomb correction. To demonstrate the impact of the Coulomb correction function and $g_{A}$ quenching, we modify these in MARLEY for the predictions in Fig 16. Details on MARLEY’s approach to the Coulomb correction function can be found in Ref. [24].
Figure 16: Predictions for the $\nu_{e}-{}^{132}$eCC cross section from MARLEY using GT matrix elements from Ref. [35] are compared with predictions from Ref. [23, 22, 21] using a different quenched $g_{A,eff}$ values and different Coulomb correction functions.

### v2

@@ VII Acknowledgments
We would like to thank the authors of Ref. [31] for providing their calculations of the GT strength used in this publication, and Steven Gardiner for his continued development and support of the MARLEY event generator. The authors gratefully acknowledge support for nEXO from the Office of Nuclear Physics within DOE’s Office of Science, and NSF in the U.S.; from NSERC, CFI, FRQNT, NRC, and the McDonald Institute (CFREF) in Canada; from IBS in Korea; and from CAS and NSFC in China. This work was supported in part by Laboratory Directed Research and Development (LDRD) programs at Brookhaven National Laboratory (BNL), Lawrence Livermore National Laboratory (LLNL), Oak Ridge National Laboratory (ORNL), Pacific Northwest National Laboratory (PNNL), and SLAC National Accelerator Laboratory. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344.
Appendix A ${}^{132}\mbox{Xe}$ inclusive cross section comparison
As a further test of MARLEY, we calculate predictions for the $\nu_{e}-{}^{132}$eCC cross section as a function of energy using GT strength from Ref. [31] and an IAS location predicted by Eq. 2. Figure 10 compares MARLEY’s predictions with existing calculations for 132Xe from Refs. [20, 22, 21]. References [20, 21] use independent theoretical calculations that include forbidden transitions and use an approach to modeling the Coulomb correction similar to that used by MARLEY. Reference [20] assumes an unquenched $g_{A}$ of $\sim$1.26. Reference  [22] uses the same theoretical GT matrix elements as we use in MARLEY, but assumes a different location for the IAS in 132Cs, an unquenched $g_{A}$ of $\sim$1.26, and uses a different approach to modeling the Coulomb correction. To demonstrate the impact of the Coulomb correction function and $g_{A}$ quenching, we modify these in MARLEY for the predictions in Fig 10. Details on MARLEY’s approach to the Coulomb correction function can be found in Ref. [23].
Figure 10: Predictions for the $\nu_{e}-{}^{132}$eCC cross section from MARLEY using GT matrix elements from Ref. [31] are compared with predictions from Refs. [22, 21, 20] using different quenched $g_{A,eff}$ values and different Coulomb correction functions.
Appendix B Individual output channel energy distributions
@@
Figs.  11, 12, and 13 show the fraction of the incident neutrino energy transferred to the various output channels for $\nu_{e}-e^{-}$, $\nu_{e}-{}^{134}$eCC, and $\nu_{e}-{}^{136}$eCC events.
Appendix C Individual transfer matrices
@@
Figs. 14,15, and16 show the transfer matrices for each of the scattering components described within the text, generated with $1\times 10^{6}$ simulated and reconstructed events with a uniform neutrino distribution from 0 to 100 MeV. These are combined with isotopic and cross section weighting in Fig. 6 of the main text.